# Some Primal-Dual Theory for Subgradient Methods for Strongly Convex Optimization

### v3

@@ Abstract
We consider (stochastic) subgradient methods for strongly convex but potentially nonsmooth non-Lipschitz optimization. We provide new equivalent dual descriptions (in the style of dual averaging) for the classic subgradient method, the proximal subgradient method, and the switching subgradient method. These equivalences enable $O(1/T)$ convergence guarantees in terms of both their classic primal gap and a not previously analyzed dual gap for strongly convex optimization. Consequently, our theory provides these classic methods with simple, optimal stopping criteria and optimality certificates at no added computational cost. Our results apply to a wide range of stepsize selections and of non-Lipschitz ill-conditioned problems where the early iterations of the subgradient method may diverge exponentially quickly (a phenomenon which, to the best of our knowledge, no prior works address). Even in the presence of such undesirable behaviors, our theory still ensures and bounds eventual convergence.
@@ 1 Introduction
where the functions $f_{s}$ are (strongly) convex but may be nonsmooth and not globally Lipschitz continuous and $r$ is convex, closed, and simple, all defined over a finite-dimensional Euclidean space $\mathcal{E}$. We will consider iteratively solving problems of this general form via a stochastic switching proximal subgradient method. This general method corresponds to the classic subgradient method when $r=0$ and $m=0$, the proximal subgradient method when $m=0$, and the switching subgradient method of [10] when $r=0$. Formal assumptions and definitions are deferred to Section 2.
@@
where the first line replaces $f_{0}$ by the maximum of its subgradient lower bounds and the second extends this to the maximum combination of such lower bounds. Any point $x\in\mathcal{E}$ provides a “primal solution” with primal gap $f_{0}(x)-p_{\star}$. Any collection of points $y_{k}$ and weights $\lambda_{k}$ provide a “dual solution”, which produces a lower bound on $f_{0}(x)$ of $m^{(T)}(x)=\frac{\sum_{k=0}^{T}\lambda_{k}\left(f_{0}(y_{k})+\langle g_{0}(y_{k}),x-y\rangle+\frac{\mu}{2}\|x-y_{k}\|_{2}^{2}\right)}{\sum_{k=0}^{T}\lambda_{k}}$ and a dual gap $p_{\star}-\inf m^{(T)}$.
The subgradient method builds a sequence of primal solutions $x_{k}$ by repeatedly moving in negative subgradient directions using a sequence of stepsizes $\alpha_{k}>0$
@@
The method of dual averaging builds dual solutions $m^{(k)}$ by repeatedly minimizing this lower bounding model (with an optional $\beta_{k}$ regularization term) and then incorporating a new subgradient into the next model from the resulting point $y_{k+1}$ with weight $\lambda_{k+1}>0$
$\begin{cases}m^{(k)}(y)=\dfrac{\sum_{i=0}^{k}\lambda_{i}\left(f_{0}(y_{i})+\langle g_{0}(y_{i}),y-y_{i},\rangle+\frac{\mu}{2}\|y-y_{i}\|_{2}^{2}\right)}{\sum_{i=0}^{k}\lambda_{i}}\\ y_{k+1}=\operatornamewithlimits{argmin}\left\{m^{(k)}(y)+\dfrac{\beta_{k}}{2\sum_{i=0}^{k}\lambda_{i}}\|y-y_{0}\|_{2}^{2}\right\}\ .\end{cases}$ (1.4)
@@
In the not strongly convex setting of $\mu=0$, Nesterov [3] showed when $\beta_{k}=\bar{\beta}>0$ is constant, these two methods are equivalent whenever $\alpha_{k}=\lambda_{k}/\bar{\beta}$. That is, they produce the same sequence of iterates $x_{k}=y_{k}$. Our Theorem 3.1 extends this equivalence to potentially strongly convex settings, showing these methods are equivalent whenever $\alpha_{k}=\lambda_{k}/(\mu\sum_{i=0}^{k}\lambda_{i}+\bar{\beta})$ and $\mu+\bar{\beta}>0$. Such results allow one to equivalently view the classic subgradient method as either iteratively building a primal solution converging down to optimal or building a lower bound converging up to optimal.
This new dual understanding enables us to develop primal-dual convergence theory for the subgradient method. We find the dual model built by the subgradient method converges up to optimal at the same rate as the iterates converge down. For example, if $\bar{\beta}=0$ and $\|g_{0}(x_{k})\|_{2}\leq M$ uniformly, a special case of our Theorem 3.2 implies for any selection of $\alpha_{k}$ and $\lambda_{k}$ with $\alpha_{k}=\lambda_{k}/(\mu\sum_{i=0}^{k}\lambda_{i})$, the subgradient method (or equivalently dual averaging) has primal gap, dual gap, and distance to optimal all converge with
@@
Given our primal-dual equivalence, our results can be equally seen as providing primal-dual convergence guarantees for dual averaging. In this sense, we improve the prior best dual averaging theory due to Deng et al. [12, Corollary 8] who showed a primal rate of $O\left(M^{2}/\mu T\right)$ when $\lambda_{k}=k+1$ and $\beta_{k}=\mu(T+2)$ is constant. Note this rate is optimal (by the example presented in [13]), meaning no faster objective gap convergence rate in terms of any of $M,\mu,T$ can be achieved.
@@ 1.1 Our Contributions
  • Dual Equivalences and Primal-Dual Convergence Theory. Our theory considers a Stochastic Switching Proximal Subgradient Method for the general problem class (1.1), which, as special cases, captures projected, proximal, and switching subgradient methods as well as gradient descent. We introduce a new dual averaging method for this general problem class, Stochastic Lagrangian Proximal Dual Averaging, which our Theorem 3.1 shows is equivalent to the stochastic switching proximal subgradient method under proper selection of primal stepsizes $\alpha_{k}$ and dual weights $\lambda_{k}$. From this equivalence, our Theorem 3.2 presents new primal-dual convergence rate guarantees for these general methods.
@@
  • Non-Lipschitz Analysis Bounds New Early Divergence Phenomena. Often, nonsmooth optimization analysis focuses on Lipschitz continuous functions. Such theory is limited to functions that asymptotically grow at most linearly. Our analysis uses a non-Lipschitz model, allowing up to quadratic growth. By doing so, our theory provides new linear primal-dual convergence guarantees for gradient descent in smooth optimization and new guarantees for minimizing a sum $f_{0}=h_{1}+h_{2}$ with smooth $h_{1}$ and nonsmooth but Lipschitz $h_{2}$, which is overall neither Lipschitz nor smooth. Numerics showcasing highly non-monotone behaviors of the subgradient method on such problems are shown in Section 4, where our theory still provides reasonably accurate predictions.
@@ 2 Preliminaries and Algorithm Definitions
First, we allow for additive composite objectives $f_{0}+r$ for any closed convex $r\colon\mathcal{E}\rightarrow\mathbb{R}\cup\{\infty\}$. We address this added term by assuming $r$ is sufficiently simple that its proximal operator $\mathrm{prox}_{\alpha_{k},r}(z):=\operatornamewithlimits{argmin}_{x}\left\{r(x)+\frac{1}{2\alpha_{k}}\|x-z\|_{2}^{2}\right\}$ can be evaluated at each iteration. For example, if $r$ is an indicator function for a closed convex constraint set, its proximal operator corresponds to projection onto that set. Since $r(x)+\frac{1}{2\alpha_{k}}\|x-z\|_{2}^{2}$ is strongly convex, it has a unique minimizer described by the following equivalence
@@
Second, we allow for $m$ strongly convex functional constraints $f_{s}(x)\leq 0$ for $s=1,\dots,m$. We address these added terms by “switching”: Given a current iterate $x$, only one function $f_{s(x)}$ will be considered that iteration. This function will be chosen as $s(x)=0$ if $x$ is feasible (that is, $f_{s}(x)\leq 0$ for all $s=1\dots m$), otherwise $s(x)$ can be chosen generically as any violated constraint $f_{s(x)}(x)>0$. Third, we allow for stochasticity in our subgradient oracles for each function, denoted by $g_{s}(x;\xi)$ such that $\mathbb{E}_{\xi}g_{s}(x;\xi)\in\partial f_{s}(x)$. Note this trivially captures deterministic methods by selecting $g_{s}(x;\xi)$ constant with respect to $\xi$.
@@
A General Primal Subgradient Method. As a primal algorithm, consider the Stochastic Switching Proximal Subgradient Method with stepsizes $\alpha_{k}>0$ and iterates $x_{k}$ defined by
$x_{k+1}=\begin{cases}\mathrm{prox}_{\alpha_{k},r}(x_{k}-\alpha_{k}g_{s(x_{k})}(x_{k};\xi_{k}))&\text{if $x_{k}$ is feasible}\\ x_{k}-\alpha_{k}g_{s(x_{k})}(x_{k};\xi_{k})&\text{otherwise.}\end{cases}$ (2.2)
for i.i.d. sampled $\xi_{k}$. When $m=0$, this is the standard (stochastic) proximal subgradient method; when $r=0$, this is the (stochastic) switching subgradient method. Note when $m>0$, only limited stochasticity can be allowed since an exact determination of feasibility is required to decide the switching variable $s(x_{k})$.
@@
$=\min_{x\in\mathcal{E}}\max_{\begin{subarray}{c}y_{0},\dots,y_{T}\in\mathcal{E}\\ \lambda_{0},\dots,\lambda_{T}\geq 0\\ \sum_{k:s(y_{k})=0}\lambda_{k}>0\\ n_{k+1}\in\partial r(y_{k+1})\end{subarray}}\mathbb{E}_{\xi}\Bigg{(}\sum_{s=0}^{m}\sum_{k:s(y_{k})=s}\lambda_{k}\left(f_{s(y_{k})}(y_{k})+\langle g_{s(y_{k})}(y_{k};\xi_{k}),x-y_{k}\rangle+\frac{\mu}{2}\|x-y_{k}\|_{2}^{2}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{k:s(y_{k})=0}\lambda_{k}\left(r(y_{k+1})+\langle n_{k+1},x-y_{k+1}\rangle\right)\Bigg{)}\frac{1}{\sum_{k:s(y_{k})=0}\lambda_{k}}$
@@
As a dual algorithm, consider the following Stochastic Lagrangian Proximal Dual Averaging with dual weights $\lambda_{k}>0$ and regularization parameters $\beta_{k}$. Its sequence of iterates $y_{k}$ based on i.i.d. sampled $\xi_{k}$ is defined as follows: At iteration $k$, construct the following (unnormalized) aggregations for each function based on the previous $k-1$ iterations as
@@
where $n_{i+1}=\frac{-1}{\lambda_{i}}(\nabla M^{(i-1)}(y_{i+1})+g_{0}(y_{i};\xi_{i})+\mu(y_{i+1}-y_{i})+\beta_{i}(y_{i+1}-y_{0}))\in\partial r(y_{i+1})$ (see Lemma 2.1 for verification of this subdifferential containment). At iteration $k=0$, these empty summations are understood to take value zero. Then, based on the switching selection $s(y_{k})$, a new weighted model is constructed as
@@
First note that when $y_{k}$ is infeasible (and so $s(y_{k})\neq 0$), $M^{(k)}=M^{(k-1)}+U^{(k)}$. From this, the result is immediate. Now consider when $y_{k}$ is feasible (and so $s(y_{k})=0$). Since $y_{k+1}$ is the unique minimizer of $M^{(k-1)}(y)+U^{k}(y)+\frac{\beta_{k}}{2}\|y-y_{0}\|_{2}^{2}$, the necessary and sufficient optimality condition ensures $y_{k+1}$ is the unique solution to
@@
and so $y_{k+1}$ is a the unique minimizer of $M^{(k)}(y)+\frac{\beta_{k}}{2}\|y-y_{0}\|_{2}^{2}$. ∎
Note whenever $s(y_{k})=0$, this step corresponds to minimizing $r$ plus a simple quadratic. This amounts to computing a proximal operator for $r$ and so is within the assumed computational oracle model. Whenever $s(y_{k})\neq 0$, this step minimizes a simple quadratic and can be done in closed form. The following easily verifiable lemma provides a simple way to maintain the minimum of the aggregate quadratic model.

### v5

@@ Abstract
We consider (stochastic) subgradient methods for strongly convex but potentially nonsmooth non-Lipschitz optimization. We provide new equivalent dual descriptions (in the style of dual averaging) for the classic subgradient method, the proximal subgradient method, and the switching subgradient method. These equivalences enable $O(1/T)$ convergence guarantees in terms of both their classic primal gap and a not previously analyzed dual gap for strongly convex optimization. Consequently, our theory provides these classic methods with simple, optimal stopping criteria and optimality certificates at no added computational cost. Our results apply to a wide range of stepsize selections and to non-Lipschitz ill-conditioned problems where the early iterations of the subgradient method may diverge exponentially quickly (a phenomenon which, to the best of our knowledge, no prior works address). Even in the presence of such undesirable behaviors, our theory still ensures and bounds eventual convergence.
@@ 1 Introduction
where the functions $f_{s}\colon\mathcal{E}\rightarrow\mathbb{R}$, $s=0,\dots,m,$ are (strongly) convex but may be nonsmooth and not globally Lipschitz continuous and $r\colon\mathcal{E}\rightarrow\mathbb{R}\cup\{\infty\}$ is convex, closed, and simple, all defined over a finite-dimensional Euclidean space $\mathcal{E}$. We will consider iteratively solving problems of this general form via a stochastic switching proximal subgradient method. This general method corresponds to the classic subgradient method when $r=0$ and $m=0$, the proximal subgradient method when $m=0$, and the switching subgradient method of [10] when $r=0$. Formal assumptions and definitions are deferred to Section 2.
@@
where the first line replaces $f_{0}$ by the maximum of its quadratic subgradient lower bounds and the second extends this to the maximum combination of such lower bounds. Any point $x\in\mathcal{E}$ provides a “primal solution” with primal gap $f_{0}(x)-p_{\star}$. Any collection of points $y_{k}$ and weights $\lambda_{k}$ provide a “dual solution”, which produces a lower bound of $m^{(T)}(x)=\frac{\sum_{k=0}^{T}\lambda_{k}\left(f_{0}(y_{k})+\langle g_{0}(y_{k}),x-y_{k}\rangle+\frac{\mu}{2}\|x-y_{k}\|_{2}^{2}\right)}{\sum_{k=0}^{T}\lambda_{k}}$ on $f_{0}(x)$ and a dual gap $p_{\star}-\inf m^{(T)}$.
The subgradient method builds a sequence of primal solutions $\{x_{k}\}$ by repeatedly moving in negative subgradient directions using a sequence of stepsizes $\alpha_{k}>0$
@@
The method of dual averaging builds dual solutions yielding models $m^{(k)}$ by repeatedly minimizing this lower bounding model (with an optional $\beta_{k}$ regularization term) and then incorporating a new subgradient $g_{0}(y_{k+1})$ into the next model from the resulting point $y_{k+1}$ with weight $\lambda_{k+1}>0$
$\begin{cases}m^{(k)}(y)=\dfrac{\sum_{i=0}^{k}\lambda_{i}\left(f_{0}(y_{i})+\langle g_{0}(y_{i}),y-y_{i}\rangle+\frac{\mu}{2}\|y-y_{i}\|_{2}^{2}\right)}{\sum_{i=0}^{k}\lambda_{i}}\\ y_{k+1}=\operatornamewithlimits{argmin}\left\{m^{(k)}(y)+\dfrac{\beta_{k}}{2\sum_{i=0}^{k}\lambda_{i}}\|y-y_{0}\|_{2}^{2}\right\}\ .\end{cases}$ (1.4)
@@
In the not strongly convex setting of $\mu=0$, Nesterov [3] showed when $\beta_{k}=\bar{\beta}>0$ is constant, these two methods are equivalent whenever $\alpha_{k}=\lambda_{k}/\bar{\beta}$. That is, they produce the same sequence of iterates $x_{k}=y_{k}$. Our Theorem 3.1 extends this equivalence to potentially strongly convex settings, showing these two methods are equivalent whenever $\alpha_{k}=\lambda_{k}/(\mu\sum_{i=0}^{k}\lambda_{i}+\bar{\beta})$ and $\mu+\bar{\beta}>0$. Such results allow one to equivalently view the classic subgradient method as either iteratively building a primal solution converging down to optimal or building a lower bound converging up to optimal.
This new dual understanding enables us to develop primal-dual convergence theory for the subgradient method. We derive convergence rate bounds ensuring the dual model built by the subgradient method converges up to optimal at at least the same rate as the primal iterates are guaranteed to converge down. For example, if $\bar{\beta}=0$ and $\|g_{0}(x_{k})\|_{2}\leq M$ uniformly, a special case of our Theorem 3.2 implies for any selection of $\alpha_{k}$ and $\lambda_{k}$ with $\alpha_{k}=\lambda_{k}/(\mu\sum_{i=0}^{k}\lambda_{i})$, the subgradient method (or equivalently dual averaging) has primal gap, dual gap, and distance to optimality all converge with
@@
Given our primal-dual equivalence, our results can be equally seen as providing primal-dual convergence guarantees for dual averaging. In this sense, we improve the prior best dual averaging theory due to Deng et al. [12, Corollary 8] who showed a primal rate of $O\left(M^{2}/\mu T\right)$ when $\lambda_{k}=k+1$ and $\beta_{k}=\mu(T+2)$ is constant. Note such rates are optimal (by the example presented in [13]), meaning no faster objective gap convergence rate in terms of any of $M,\mu,T$ can be achieved.
@@ 1.1 Our Contributions
  • Dual Equivalences and Primal-Dual Convergence Theory. Our theory considers a Stochastic Switching Proximal Subgradient Method for the general problem class (1.1), which, as special cases, captures projected, proximal, and switching subgradient methods as well as gradient descent. We introduce a new dual averaging method for this general problem class, Stochastic Lagrangian Proximal Dual Averaging, which our Theorem 3.1 shows is equivalent to the stochastic switching proximal subgradient method under proper selection of primal stepsizes $\alpha_{k}$ and dual weights $\lambda_{k}$ and $\beta_{k}$. From this equivalence, our Theorem 3.2 presents new primal-dual convergence rate guarantees for these general methods.
@@
  • New Non-Lipschitz Analysis Bounds for Early Divergence Phenomena. Often, nonsmooth optimization analysis focuses on Lipschitz continuous functions. Such theory is limited to functions that asymptotically grow at most linearly. Our analysis uses a non-Lipschitz model, allowing up to quadratic growth. By doing so, our theory provides new linear primal-dual convergence guarantees for gradient descent in smooth optimization and new guarantees for minimizing a sum $f_{0}=h_{1}+h_{2}$ with smooth $h_{1}$ and nonsmooth but Lipschitz $h_{2}$, which is overall neither Lipschitz nor smooth. Numerics showcasing highly non-monotone behaviors of subgradient methods on such problems are shown in Section 4, where our theory still provides reasonably accurate predictions.
@@ 2 Preliminaries and Algorithm Definitions
First, we allow for additive composite objectives $f_{0}+r$ for any proper closed convex $r\colon\mathcal{E}\rightarrow\mathbb{R}\cup\{\infty\}$. We address this added term by assuming $r$ is sufficiently simple that its proximal operator $\mathrm{prox}_{\alpha_{k},r}(z):=\operatornamewithlimits{argmin}_{x}\left\{r(x)+\frac{1}{2\alpha_{k}}\|x-z\|_{2}^{2}\right\}$ can be evaluated at each iteration. For example, if $r$ is an indicator function for a closed convex constraint set, its proximal operator corresponds to projection onto that set. Since $r(x)+\frac{1}{2\alpha_{k}}\|x-z\|_{2}^{2}$ is strongly convex, it has a unique minimizer described by the following equivalence
@@
Second, we allow for $m$ strongly convex functional constraints $f_{s}(x)\leq 0$ for $s=1,\dots,m$. We address these added terms by “switching”: Given a current iterate $x$, only one function $f_{s(x)}$ will be considered in that iteration. This function will be chosen as $s(x)=0$ if $x$ is feasible (that is, $f_{s}(x)\leq 0$ for all $s=1\dots m$), otherwise $s(x)$ can be chosen generically as any violated constraint $f_{s(x)}(x)>0$. Third, we allow for stochasticity in our subgradient oracles for each function, denoted by $g_{s}(x;\xi)$ such that $\mathbb{E}_{\xi}g_{s}(x;\xi)\in\partial f_{s}(x)$. Note this trivially captures deterministic methods by selecting $g_{s}(x;\xi)$ constant with respect to $\xi$.
@@
A General Primal Subgradient Method. As a primal algorithm, consider the Stochastic Switching Proximal Subgradient Method with stepsizes $\alpha_{k}>0$ and sequence of iterates $\{x_{k}\}$ defined by
$x_{k+1}=\begin{cases}\mathrm{prox}_{\alpha_{k},r}(x_{k}-\alpha_{k}g_{0}(x_{k};\xi_{k}))&\text{if $x_{k}$ is feasible}\\ x_{k}-\alpha_{k}g_{s(x_{k})}(x_{k};\xi_{k})&\text{otherwise.}\end{cases}$ (2.2)
for i.i.d. sampled $\xi_{k}$. Throughout, we always assume $\alpha_{k}>0$, strictly. When $m=0$, this is the standard (stochastic) proximal subgradient method; when $r=0$, this is the (stochastic) switching subgradient method. Note when $m>0$, only limited stochasticity can be allowed since an exact determination of feasibility is required to decide the switching variable $s(x_{k})$.
@@
$=\min_{x\in\mathcal{E}}\max_{\begin{subarray}{c}y_{0},\dots,y_{T}\in\mathcal{E}\\ \lambda_{0},\dots,\lambda_{T-1}\geq 0\\ \sum_{k<T\colon s(y_{k})=0}\lambda_{k}>0\\ n_{k+1}\in\partial r(y_{k+1})\end{subarray}}\mathbb{E}_{\xi}\Bigg{(}\sum_{s=0}^{m}\sum_{\begin{subarray}{c}k<T\\ s(y_{k})=s\end{subarray}}\lambda_{k}\left(f_{s(y_{k})}(y_{k})+\langle g_{s(y_{k})}(y_{k};\xi_{k}),x-y_{k}\rangle+\frac{\mu}{2}\|x-y_{k}\|_{2}^{2}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{\begin{subarray}{c}k<T\\ s(y_{k})=0\end{subarray}}\lambda_{k}\left(r(y_{k+1})+\langle n_{k+1},x-y_{k+1}\rangle\right)\Bigg{)}\frac{1}{\sum_{k<T\colon s(y_{k})=0}\lambda_{k}}$
@@
As a dual algorithm, consider the following Stochastic Lagrangian Proximal Dual Averaging with dual weights $\lambda_{k}>0$ and regularization parameters $\beta_{k}\geq 0$. Throughout, we always assume $\lambda_{k}>0$, strictly. Its sequence of iterates $\{y_{k}\}$ based on i.i.d. sampled $\xi_{k}$ is defined as follows: At iteration $k$, construct the following (unnormalized) aggregations for each function based on the previous $k-1$ iterations as
@@
where $n_{i+1}=\frac{-1}{\lambda_{i}}(\nabla M^{(i-1)}(y_{i+1})+\lambda_{i}(g_{0}(y_{i};\xi_{i})+\mu(y_{i+1}-y_{i}))+\beta_{i}(y_{i+1}-y_{0}))\in\partial r(y_{i+1})$ (see Lemma 2.1 for verification of this subdifferential containment). At iteration $k=0$, these empty summations are understood to take value zero. Then, based on the switching selection $s(y_{k})$, a new weighted model is constructed as
@@
First note that when $y_{k}$ is infeasible (and so $s(y_{k})\neq 0$), $M^{(k)}=M^{(k-1)}+U^{(k)}$. From this, the result is immediate. Now consider when $y_{k}$ is feasible (and so $s(y_{k})=0$). Since $y_{k+1}$ is the unique minimizer of $M^{(k-1)}(y)+U^{(k)}(y)+\frac{\beta_{k}}{2}\|y-y_{0}\|_{2}^{2}$, the necessary and sufficient optimality condition ensures $y_{k+1}$ is the unique solution to
@@
and so $y_{k+1}$ is a the unique minimizer of $M^{(k)}(y)+\frac{\beta_{k}}{2}\|y-y_{0}\|_{2}^{2}$.
Note whenever $s(y_{k})=0$, the step (2.3) corresponds to minimizing $r$ plus a simple quadratic. This amounts to computing a proximal operator for $r$ and so is within the assumed computational oracle model. Whenever $s(y_{k})\neq 0$, the step (2.3) minimizes a simple quadratic and can be done in closed form. The following easily verifiable lemmas provide a simple way to maintain the minimum of the aggregate quadratic model.
Lemma 2.2.
Any quadratic function of the form $Q(z)=c+\langle d,z\rangle+\frac{b}{2}\|z-\hat{z}\|^{2}_{2}$ with $b>0$ is equal to $Q(z)=\min Q+\frac{b}{2}\|z-\operatornamewithlimits{argmin}Q\|^{2}_{2}$.